OpenStudy (anonymous):

find dy/dx by implicit differentiation y^4+xy^2=5+x^3y

OpenStudy (anonymous):

Implicit differentiation just means that you need to find a dy/dx from the equation, which is also seen as y'. Knowing this, if we differentiate the enter problem we should be able to find dy/dx afterwards by using algebra.

OpenStudy (valpey):

Is that:\[y^4+xy^2=5+x^3y\]or\[y^4+xy^2=5+x^{3y}\]

OpenStudy (anonymous):

first one

OpenStudy (kainui):

Since you don't know y, really when you're differentiating it you use the chain rule. That's the only thing you do in implicit differentiation.

OpenStudy (anonymous):

can you walk me through that?

OpenStudy (anonymous):

When you finding the derivative to respect to y, and you differentiate two sides of an equation, each x would differentiate normally. Ex ) x^2 = 2x. The tricky part is finding what the y's turn into.

OpenStudy (valpey):

So do you get what @Coda is saying? \[\partial (x^2y^2)=2xy^2\partial x+x^22y\partial y\]

OpenStudy (anonymous):

not really

OpenStudy (kainui):

So when you see something like 5y^2=3x^2 all you are really doing is saying (10y)dy/dx=6x*1 See on the left side you did the derivative as normal, the outside stuff, then multiplied by the derivative of the inside. The derivative of y is just dy/dx!

OpenStudy (valpey):

Or to back up one step with @Kainui \[5y^2=3x^2\]\[10y \ dy=6x\ dx\]

OpenStudy (valpey):

Don't try to reconcile the dy/dx until the end. First just take the derivatives of both sides.

OpenStudy (valpey):

Or of each term as if each term were a separate derivative.

OpenStudy (valpey):

So let's start with the first term y^4. What is the derivative of y^4?

OpenStudy (kainui):

@Coda I didn't say anything incorrect, I just gave an example. Also, remember that you're using the product rule and the chain rule at the same time, so be careful! The chain rule just means one of the products times the derivative of one plus the derivative of the other times the other kept constant.

OpenStudy (anonymous):

Sorry @Kainui, I forgot to add my commas or they did not register.

OpenStudy (lgbasallote):

@jeanniebean what dont you get here?

OpenStudy (anonymous):

umm i just need an example with this problem with the two people trying to help me i got confused and just have alot of stuff written down can we start at the beginning and go step by step?

OpenStudy (lgbasallote):

\[y^4 + xy^2 = 5 + x^3 y\] what is the derivative of y^4?

OpenStudy (anonymous):

4y^3

terenzreignz (terenzreignz):

That's almost right, but we assume for the moment that y is a function of x, so, after you're done differentiating y as you would x, you affix a dy/dx :)

OpenStudy (lgbasallote):

right but dont forget y'

OpenStudy (anonymous):

so 4dy/dx^3?

OpenStudy (lgbasallote):

no.. \(4y^3 y'\)

OpenStudy (anonymous):

oh ok

terenzreignz (terenzreignz):

No, just differentiate it as you would x,which is what you did... 4y^3 and then when you're done, multiply the whole thing with dy/dx, or y'

OpenStudy (lgbasallote):

next what is the derivative of xy^2? use product rule here

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x(2y*y')+1(y^2)

terenzreignz (terenzreignz):

That's right :) I take it you can do the rest?

OpenStudy (anonymous):

what do i do after i get 4y^3(y')+x(2y*y')+1(y^2)= 0+x^3(y')=3x^2(y)

OpenStudy (anonymous):

put y' on one side and remaining on the other side of equal to..

OpenStudy (anonymous):

how do i do that with it being so many?

OpenStudy (lgbasallote):

what are the terms with y'?

OpenStudy (anonymous):

4y^3(y')+x(2y*y')=x^3(y')

OpenStudy (anonymous):

i guess u left out y^2 take common from where u cn...result according to me is \[-y ^{2}/(4y ^{3}+2xy-x ^{3)}\]

OpenStudy (anonymous):

what is that the result for?

OpenStudy (anonymous):

that is y' or dy/dx

OpenStudy (anonymous):

oh so like thats the answer?

OpenStudy (anonymous):

thats all you have to do?

OpenStudy (anonymous):

according to me..yes!

OpenStudy (anonymous):

oh ok well can you help me with integrals?

OpenStudy (anonymous):

i cn try integrals...

OpenStudy (anonymous):

ok : \[\int\limits_{?}^{?} (2x ^{2}+ 5\sec ^{2}x)dx=\]

OpenStudy (anonymous):

there is nothing where the question marks are

OpenStudy (anonymous):

i thnk...answer is \[x ^{4}/2 + 5 tanx\]

OpenStudy (anonymous):

can you write out how you got to that?

OpenStudy (anonymous):

yeah sure..integral of any nx^y is (nx^(y+1))/y+1 and integral of sec^2 x is tan x.... so...integral of 2x^3 should hv been \[(2x ^{3+1})/(3+1)\]

OpenStudy (anonymous):

oh and then you just get the answer?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

i thnk...!

OpenStudy (anonymous):

good enough for me if i give you another one can you try it and just write out the steps that you took to get there too?

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

\[\int\limits_{1}^{2}(1-2x ^{2}+x ^{5})/x ^{5} dx=\] it is the part in parentheses over the x^5 sorry i dont know how to make a fraction

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

umm...(1-2x^3+x^5)/x^5 is the fraction??

OpenStudy (anonymous):

yeah like the part in parentheses over the x^5

OpenStudy (anonymous):

ohk...w8!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i htnk i ll do it wrong...i m yet to do my integration classes in school...

OpenStudy (anonymous):

oh ok